The sum of two numbers is 108 and their difference is 16. Find each of the numbers.
Please show work.
If you can find an answer key that would be great.

Found 2 solutions by JulietG, Edwin McCravy:
Answer by JulietG(1812)

Two numbers, right? Not "a number"?.
.
Let A be the first number
Let B be the second number
Here's what we know:
A + B = 108
A - B = 16
Add these together
2A = 124 (the +B and -B have canceled each other out)
Divide by 2
A = 62
.
If A is 62, then B is 108-62, or 46.
Let's plug it in to the second equation to prove it.
A - B = 16
62 - 46 = 16
Success!

Answer by Edwin McCravy(20059)

 x + y = 108   <-- that says "their sum is 108"
 x - y =  16   <-- that says "their difference is 16"   
------------

Add the two equations term by term.
Then solve for x.
Then substitute the value you get for x into
either one of those equations and solve for y.

Edwin